# Effective Dynamics of Spherically Symmetric Static Spacetime

### v1

@@ I Introduction
Gravity universally interacts with all types of matter and fields, inevitably intertwining their dynamics in a complex non-linear dance. Spherically symmetric static space-times provide the most basic arena for investigating such universal dynamics of gravity. Due to its role as universal equilibrium (non-rotating) geometries, this symmetric-reduced configuration plays, despite its simplicity, a significant role in various contexts, among which the analysis of the internal structure of self-gravitating compact objects (e.g. Tolman:1939jz ; Oppenheimer:1939ne ; Buchdahl:1959zz ; Cardoso:2019rvt ), the exploration of quantum black holes (e.g. Kuchar:1994zk ; Ashtekar:2005qt ; Haggard:2014rza ) and more generally of the quantum gravity regime (e.g. Thiemann:1992jj ; Bojowald:2005cb ). In this context, we investigate how to model the feedback of classical and quantum matter fields, as well as modified gravitational dynamics, on spherically symmetric static spacetime through simple consistent action principles and effective Hamiltonians, and how these modify the structure of black hole solutions.
@@
More precisely, we consider the following metric ansatz for spherically-symmetric static space-time111 A slight change of gauge is the line element ansatz: $\mathrm{d}s^{2}=-F(R)\mathrm{d}t^{2}+F(R)^{-1}\mathrm{d}R^{2}+r(R)^{2}\mathrm{d}\Omega^{2}\,,$ where $F$ and $r$ are now functions of the new radial coordinate $R$. The correspondance is highly non-linear and explicitly given by $\mathrm{d}_{R}r=b(r)$ and $F(R)=f(r(R))b(r(R))^{2}$. :
@@
where the metric components are solely functions of the radial coordinate $r$. The field $b(r)$ is the lapse up to the factor $f(r)$. It is more exactly the area element of the 2D surface normal to the spatial sphere with radius $r$. Further introducing the field $a(r)$ such that $f(r)=1-a(r)/{r}$, we recognize $m(r)=a(r)/2G$ as the Misner-Sharp mass contained in the ball of radius $r$. It is a locally conserved energy (even in dynamical cases Kodama ) with contributions from intrinsic mass and gravitational energy, which agrees with the ADM energy in the $r\to\infty$ limit for asymptotically flat spacetimes MTW ; Hayward:1994bu .
@@
[…]
@@
such that $H_{\mathrm{eff}}$ is linear in $b$. Here, $\rho_{\mathrm{eff}}\equiv-4\pi r^{2}{\mathcal{T}}^{{\rm eff}}{}^{t}{}_{t}$ is the 1D energy density, and the coefficient functions $c_{n}(r)$ characterize the system of interest. This ansatz is consistent with the logic that the lapse factor $b$ plays the role of the Lagrange multiplier that enforces general relativity’s Hamiltonian constraint generating diffeomorphisms in the radial direction. In section V, we show that the feedback of fluids, more particularly the cases of conformal fluid and causal limit fluid, on the geometry can indeed be encoded in such a Hamiltonian.
@@
This formalism, although very simple in its conceptual framework and practical implementation, seems versatile enough to explore the dynamics of the geometry consistently coupled with renormalized quantum fields with gravitational feedback, at least for spherically symmetric static space-times. Indeed it allows for a systematic exploration of modified Schwarzschild black holes, possibly singularity-free, together with their matter content on both sides of the Schwarzschild radius.
@@ II.1 Pure gravity
with $f(r)=1-a(r)/r$. The function $a(r)$ directly gives the Misner-Sharp mass $m(r)=a(r)/2G$ of the spatial ball of radius $r$.
@@
where $A(r)\equiv 4\pi r^{2}$, $\kappa(r)=\sqrt{-g_{tt}(r)}\alpha(r)$ is the surface gravity at $r$, and $\alpha(r)\equiv|\alpha_{\mu}\alpha^{\mu}|^{1/2}$ is the proper acceleration. This agrees with the on-shell Noether charge for time diffeomorphisms (or equivalently, the Komar charge for the Killing vector $\partial_{t}$) Compere . Moreover, it is well-knwown that it is related to the entropy in black hole thermodynamics222 More precisely, by using a path-integral representation of the density of states $\nu$, the entropy can be obtained as $\mathcal{S}=\ln\nu\approx\frac{1}{\hbar}l_{BH}\mathcal{Q}_{t}(a_{0})=\frac{A(a_{0})}{4\hbar G}$, where $a_{0}$ is the Schwarzschild radius, $l_{BH}=\hbar\beta_{BH}=\frac{2\pi}{\kappa(a_{0})}$ is the periodic (Euclidean) time IyerWald . It can also obtained as the adiabatic-invariant and integrable Hamiltonian of (dynamical) horizons for “thermal time flow” generated by $\hbar\beta_{BH}\partial_{t}$ PY . .
@@ II.2 Effective Hamiltonian
To go beyond pure gravity and general relativity, a natural direction of investigation is to simply explore the possibility of a non-vanishing Hamiltonian for our canonical pair $(a(r),b(r))$. Indeed, under the two assumptions that we keep focussing on the static spherically symmetric equilibirum configurations (and thus do not look into out-of-equilibrium dynamics), and that the relevant physical degrees of freedom describing spherically symmetric spacetime remain333 Let us point out that this assumption of keeping the canonical pair $(a(r),b(r))$ excludes higher derivative terms in the Lagrangian, since those would generate extra degrees of freedom, formalized as higher momenta in the Hamiltonian formulation (see e.g. Woodard:2015zca ). In particular, introducing higher curvature terms in the Einstein-Hilbert action, such as $R^{2}$ or $R_{\mu\nu}R^{\mu\nu}$, would enrich our reduced gravitational phase space with extra canonical pairs. the mass $a$ and lapse factor $b$, the only way to generate non-trivial profiles $(a(r),b(r))$ (in terms of the radial coordinate) is to introduce a non-vanishing Hamiltonian, as long as the studied effects can be modeled by a Lagrangian or Hamiltonian action principle.
@@
    As the first equation is a decoupled differential equation solely on the mass $a(r)$, this system is much simpler to solve than the type $(-1)$. Once the evolution of the mass is integrated, one can plug it into the $b$-equation and obtain the lapse evolution.
@@ III.2 Massless scalar field
where the scalar field momentum $p_{\phi}$ now plays the role of a coupling constant in front of the effective Hamiltonian444 This is a local function of $r$, obtained by integrating out the scalar field, but it contains non-local information in the sense that the value of the constant $p_{\phi}$ is determined by a boundary condition. ,
@@ III.4 Non-linear Electromagnetic Theory
$S_{\mathrm{NL}}$ $=\int\mathrm{d}x^{4}\,\sqrt{-g}\,{\cal L}(\hat{F})=4\pi l_{0}\int\mathrm{d}r\,r^{2}b{\cal L}(\hat{F})$ (53)
where ${\cal L}(\hat{F})$ is an arbitrary function of the field strength $\hat{F}$ given earlier in (44).
@@
$S_{\mathrm{tot}}=\int\mathrm{d}r\,\left(\frac{b\dot{a}}{2G}+p_{A}\dot{A}-H_{EM}\right)\,,$ (54)
@@
$p_{A}=\frac{4\pi r^{2}}{b}\dot{A}\partial_{\hat{F}}{\cal L}\,,\quad H_{EM}=b\left[\frac{p_{A}^{2}}{4\pi r^{2}\partial_{\hat{F}}{\cal L}}-4\pi r^{2}{\cal L}\right]\,.$
@@
$S_{\mathrm{eff}}^{(EM)}$ $=\int\mathrm{d}r\,\left(\frac{b\dot{a}}{2G}-H_{\mathrm{eff}}^{(EM)}\right)$ (55)
$\textrm{with}\quad H_{\mathrm{eff}}^{(EM)}[a,b,r]=b\left[\frac{p_{A}^{2}}{4\pi r^{2}\partial_{\hat{F}}{\cal L}}-4\pi r^{2}{\cal L}\right]\,,$
where the field momentum $p_{A}$ plays the role of a coupling constant for the geometry. This is always an effective Hamiltonian of type $(+1)$ as for the standard Maxwell field. Moreover, since the field strength $\hat{F}$ does not depend on $a$, the effective Hamiltonian never also depends directly on the metric component $a$, so the lapse factor is always constant on classical solutions, $\dot{b}=0$, and only the dynamics of the mass $a$ is affected by the non-linear extension of the EM field.
Now that we have found that the simplest matter fields explicitly provide examples of effective Hamiltonian for the geometrical sector, particularly of types $(-1)$ and $(+1)$, we can turn to the general constraints of the effective Hamiltonian set by its compatibility with the Einstein equations and, in particular, with fluid matter.
@@ IV.1 Compatibility with Einstein equations
where $\rho\equiv 4\pi r^{2}(-{\cal T}{}^{t}{}_{t})$ and $p_{r}\equiv 4\pi r^{2}{\cal T}{}^{r}{}_{r}$ are respectively the energy density and radial pressure averaged over the sphere of radius $r$. Comparing to the equations of motion (6) for the effective Hamiltonian implies that
@@ IV.2 Linear Equations of State
where $\eta$ is a constant parameter characterizing the fluid. This can be applied to both a locally isotropic fluid (${\cal T}{}^{r}{}_{r}={\cal T}{}^{\theta}{}_{\theta}$ Nilsson:2000zf ) and an anisotropic material (${\cal T}{}^{r}{}_{r}\neq{\cal T}{}^{\theta}{}_{\theta}$). Requiring the dominant energy condition $p_{\mathrm{eff}}\leq\rho_{\mathrm{eff}}$, and the non-negativity of $p_{\mathrm{eff}}$ and $\rho_{\mathrm{eff}}$, imply that $\eta$ must satisfy
$1\leq\eta\leq 2\,.$ (62)
@@
without any specific constraint on the way it depends on the radial coordinate $r$. This allows to recover in a straightforward fashion the various examples of matter explored in the previous section. Indeed, for the cosmological constant or a Maxwell field, one has ${\cal T}^{t}{}_{t}={\cal T}^{r}{}_{r}$, corresponding to $\eta=\infty$, implying that the effective Hamiltonian should not depend on the field $a$ but solely on $b$. This is indeed consistent with what was derived earlier in (27) and (48). As for the massless scalar field, one has $-{\cal T}^{t}{}_{t}={\cal T}^{r}{}_{r}$, indicating $\eta=1$ and thus implying that the effective Hamiltonian should be a function of $c=b(1-a/r)$. This is perfectly consistent with our earlier derivation of the effective Hamiltonian (34) induced by a massless scalar field, which we recall here:
@@ V.1 Polynomial ansatz in powers of mass
$\left|\begin{array}[]{l}\rho_{\mathrm{eff}}[a,r]=c_{0}(r)\,,\vspace*{1mm}\\ c_{n}(r)=0\,\,{\rm{for}}\,\,n\geq 1\,,\end{array}\right.\quad\Rightarrow\quad\begin{array}[]{l}\dot{a}=c_{0}\,,\vspace*{1mm}\\ \dot{b}=0\,,\end{array}$ (68)
@@ V.2 Effective Hamiltonians for Fluid Matter
To illustrate the scope of the presented framework, let us construct the effective Hamiltonian for self-gravitating isotropic fluids, whose effective action is difficult to construct directly from a fundamental matter Lagrangian.
@@
Let us look for an effective Hamiltonian of type $(+1)$, thus an effective energy density $\rho_{\mathrm{eff}}[a,r]$, that fits with this evolution $(a(r),b(r))$. Let us assume a quadratic truncation with
@@
$2G\rho_{\mathrm{eff}}=\frac{9}{14}+\frac{9}{147}c_{2}r^{2}-\left(\frac{1}{2r}+\frac{3}{7}c_{2}r\right)a+\frac{1}{2}c_{2}a^{2}\,,$ (77)
where $c_{2}(r)$ is an arbitrary function. Let’s not forget (from (60)) that we still need to impose $\frac{\partial\rho_{\mathrm{eff}}(a,r)}{\partial a}<0$ to respect the positive energy condition. This constraint is automatically satisfied for the simplest choice $c_{2}(r)=0$, which reads:
@@
Next, we consider a causal-limit fluid. It is defined by the isotropy condition $T^{r}{}_{r}=T^{\theta}{}_{\theta}\equiv p_{3d}$ and the equation of state $p_{3d}=\rho_{3d}(\equiv-T^{t}{}_{t})$, which saturates the dominant energy condition and corresponds to the limiting situation in which causal propagation is established Zeldovich:1961sbr . As shown in e.g. Appendix A of Y2 , these two conditions determine the metric of the spherically-symmetric static configuration with size $\ell$:
@@
where $f(r)=1-{a(r)}/{r}$ is the usual metric factor entering the spherically-symmetric ansatz (3). We do not have a simple argument for obtaining this surprisingly simple effective Hamiltonian. Nevertheless, we point out that $bf$ is simply $\sqrt{{-g_{tt}}/{g_{rr}}}$, that is, the speed of light for radial light-like geodesics. This might be more that a mere coincidence for the spherically-symmetric causal-limit fluid, which has no characteristic properties except for the limiting behavior $\rho_{3d}=p_{3d}$, but we postpone the investigation of a deeper origin of this effective Hamiltonian to future work.

### v2

@@ I Introduction
Gravity universally interacts with all types of matter and fields, inevitably intertwining their dynamics in a complex non-linear dance. Spherically symmetric static space-times provide the most basic arena for investigating such universal dynamics of gravity. Due to its role as universal equilibrium (non-rotating) geometries, this symmetric-reduced configuration plays, despite its simplicity, a significant role in various contexts, among which the analysis of the internal structure of self-gravitating compact objects (e.g. Tolman:1939jz ; Oppenheimer:1939ne ; Buchdahl:1959zz ; Cardoso:2019rvt ), the exploration of quantum black holes (e.g. Kuchar:1994zk ; Ashtekar:2005qt ; Haggard:2014rza ) and more generally of the quantum gravity regime (e.g. Thiemann:1992jj ; Bojowald:2005cb ). In particular, consideration of a radial Hamiltonian for spherically symmetric spacetime is receiving attention recently (e.g.Bodendorfer:2019cyv ; Perez:2023jrq ; BenAchour:2023dgj ; MenaMarugan:2024 ; Koch:2025yuz ). In this context, we investigate how to model the feedback of classical and quantum matter fields, as well as modified gravitational dynamics, on spherically symmetric static spacetime through simple consistent action principles and effective Hamiltonians, and how these modify the structure of black hole solutions.
@@
More precisely, we consider the following metric ansatz for spherically-symmetric static space-time [^1] In other works (e.g. Bodendorfer:2019cyv ; BenAchour:2023dgj ; MenaMarugan:2024 ), a different form of the line element ansatz is used: $\mathrm{d}s^{2}=-F(R)\mathrm{d}t^{2}+F(R)^{-1}c(R)^{2}\mathrm{d}R^{2}+r(R)^{2}\mathrm{d}\Omega^{2}.$ Here, addition to the lapse $c(R)$ and a function $F(R)$, $r(R)$ is a function of a new radial coordinate $R$. The number of degrees of freedom seems different from ours, but after fixing the gauge, they are the same. Indeed, setting $c(R)=1$ is a gauge fixing, and the correspondence is given by $\mathrm{d}_{R}r=b(r)$ and $F(R)=f(r(R))b(r(R))^{2}$. Note that, at least, one of the key results, (4), holds even in a full covariant formalism PY .:
@@
where the metric components are solely functions of the radial coordinate $r$. The field $b(r)$ is the lapse up to the factor $f(r)$. It is more exactly the area element of the 2D surface normal to the spatial sphere with radius $r$. Further introducing the field $a(r)$ such that f(r)=1a(r)/rf(r)=1-a(r)/{r}, we recognize $m(r)=a(r)/2G$ as the Misner-Sharp mass contained in the ball of radius $r$. It is a locally conserved energy (even in dynamical cases Kodama ) with contributions from intrinsic mass and gravitational energy, which agrees with the ADM energy in the $r\to\infty$ limit for asymptotically flat spacetimes MTW ; Hayward:1994bu .
@@
$S_{\mathrm{g}}[a,b]=\frac{1}{2G}\int^{r_{2}}_{r_{1}}drb(r)\dot{a}(r),\quad\textrm{with}\,\,\dot{a}\equiv\frac{\mathrm{d}a}{\mathrm{d}r},$ (4)
@@
such that $H_{\mathrm{eff}}$ is linear in $b$. Here, $\rho_{\mathrm{eff}}\equiv-4\pi r^{2}{\mathcal{T}}^{{\rm eff}}{}^{t}{}_{t}$ is the 1D energy density, and the coefficient functions $c_{n}(r)$ characterize the system of interest. This ansatz is consistent with the logic that the lapse factor $b$ plays the role of the Lagrange multiplier that enforces general relativity’s Hamiltonian constraint generating diffeomorphisms in the radial direction, and that the nonlinear self-interaction of gravity can be modeled by higher powers of mass $a$. In section V, we show that the feedback of fluids, more particularly the cases of conformal fluid and causal limit fluid, on the geometry can indeed be encoded in such a Hamiltonian.
@@
This formalism, although very simple in its conceptual framework and practical implementation, seems versatile enough to explore the dynamics of the geometry consistently coupled with renormalized quantum fields with gravitational feedback and the dynamics beyond general relativity, at least for spherically symmetric static space-times. Indeed it allows for a systematic exploration of modified Schwarzschild black holes, possibly singularity-free, together with their matter content on both sides of the Schwarzschild radius.
@@ II.1 Pure gravity
with f(r)=1a(r)/rf(r)=1-a(r)/r. The function $a(r)$ directly gives the Misner-Sharp mass $m(r)=a(r)/2G$ of the spatial ball of radius $r$.
@@
where $A(r)\equiv 4\pi r^{2}$, $\kappa(r)=\sqrt{-g_{tt}(r)}\alpha(r)$ is the surface gravity at $r$, and $\alpha(r)\equiv|\alpha_{\mu}\alpha^{\mu}|^{1/2}$ is the proper acceleration. This agrees with the on-shell Noether charge for time diffeomorphisms (or equivalently, the Komar charge for the Killing vector $\partial_{t}$) Compere . Moreover, it is well-knwown that it is related to the entropy in black hole thermodynamics [^2] More precisely, by using a path-integral representation of the density of states $\nu$, the entropy can be obtained as $\mathcal{S}=\ln\nu\approx\frac{1}{\hbar}l_{BH}\mathcal{Q}_{t}(a_{0})=\frac{A(a_{0})}{4\hbar G}$, where $a_{0}$ is the Schwarzschild radius, $l_{BH}=\hbar\beta_{BH}=\frac{2\pi}{\kappa(a_{0})}$ is the periodic (Euclidean) time IyerWald . It can also obtained as the adiabatic-invariant and integrable Hamiltonian of (dynamical) horizons for “thermal time flow” generated by $\hbar\beta_{BH}\partial_{t}$ PY . .
@@ II.2 Effective Hamiltonian
To go beyond pure gravity and general relativity, a natural direction of investigation is to simply explore the possibility of a non-vanishing Hamiltonian for our canonical pair $(a(r),b(r))$. Indeed, under the two assumptions that we keep focussing on the static spherically symmetric equilibirum configurations (and thus do not look into out-of-equilibrium dynamics), and that the relevant physical degrees of freedom describing spherically symmetric spacetime remain [^3] Let us point out that this assumption of keeping the canonical pair $(a(r),b(r))$ excludes higher derivative terms in the Lagrangian, since those would generate extra degrees of freedom, formalized as higher momenta in the Hamiltonian formulation (see e.g. Woodard:2015zca ). In particular, introducing higher curvature terms in the Einstein-Hilbert action, such as $R^{2}$ or $R_{\mu\nu}R^{\mu\nu}$, would enrich our reduced gravitational phase space with extra canonical pairs. the mass $a$ and lapse factor $b$, the only way to generate non-trivial profiles $(a(r),b(r))$ (in terms of the radial coordinate) is to introduce a non-vanishing Hamiltonian, as long as the studied effects can be modeled by a Lagrangian or Hamiltonian action principle.
@@
    As the first equation is a decoupled differential equation solely on the mass $a(r)$, this system is much simpler to solve than the type (1)(-1). Once the evolution of the mass is integrated, one can plug it into the $b$-equation and obtain the lapse evolution.
@@ III.2 Massless scalar field
where the scalar field momentum $p_{\phi}$ now plays the role of a coupling constant in front of the effective Hamiltonian [^4] This is a local function of $r$, obtained by integrating out the scalar field, but it contains non-local information in the sense that the value of the constant $p_{\phi}$ is determined by a boundary condition. ,
@@ III.4 Non-linear Electromagnetic Theory
[…]
@@
$S_{\mathrm{tot}}=\int\mathrm{d}r\,\left(\frac{b\dot{a}}{2G}+p_{A}\dot{A}-H_{\mathrm{NL}}\right)\,$, (54)
@@
$p_{A}=\frac{4\pi r^{2}}{b}\dot{A}\partial_{\hat{F}}{\cal L}\,,\quad H_{\mathrm{NL}}=b\left[\frac{p_{A}^{2}}{4\pi r^{2}\partial_{\hat{F}}{\cal L}}-4\pi r^{2}{\cal L}\right]\,.$
@@
$S_{\mathrm{eff}}^{(NL)}$ $=\int\mathrm{d}r\,\left(\frac{b\dot{a}}{2G}-H_{\mathrm{eff}}^{(NL)}\right)$ (55)
$\textrm{with}\quad H_{\mathrm{eff}}^{(NL)}[a,b,r]=b\left[\frac{p_{A}^{2}}{4\pi r^{2}\partial_{\hat{F}}{\cal L}}-4\pi r^{2}{\cal L}\right]\,$,
where the field momentum $p_{A}$ plays the role of a coupling constant for the geometry although the $p_{A}$-dependence is intricate generically. This is always an effective Hamiltonian of type $(+1)$ as for the standard Maxwell field. Moreover, since the field strength $\hat{F}$ does not depend on $a$, the effective Hamiltonian never also depends directly on the metric component $a$, so the lapse factor is always constant on classical solutions, $\dot{b}=0$, and only the dynamics of the mass $a$ is affected by the non-linear extension of the EM field.
Now that we have found that the simplest matter fields explicitly provide examples of effective Hamiltonian for the geometrical sector, particularly of types (1)(-1) and $(+1)$, we can turn to the general constraints of the effective Hamiltonian set by its compatibility with the Einstein equations and, in particular, with fluid matter.
@@ IV.1 Compatibility with Einstein equations
[…]
@@ IV.2 Linear Equations of State
where $\eta$ is a constant parameter characterizing the fluid. This can be applied to both a locally isotropic fluid (${\cal T}{}^{r}{}_{r}={\cal T}{}^{\theta}{}_{\theta}$ Nilsson:2000zf ) and an anisotropic material (${\cal T}{}^{r}{}_{r}\neq{\cal T}{}^{\theta}{}_{\theta}$). Requiring the dominant energy condition $p_{\mathrm{eff}}\leq\rho_{\mathrm{eff}}$, and the positivity of $p_{\mathrm{eff}}$ and $\rho_{\mathrm{eff}}$, imply that $\eta$ must satisfy
$1\leq\eta<2\,.$ (62)
@@
without any specific constraint on the way it depends on the radial coordinate $r$. This allows to recover in a straightforward fashion the various examples of matter explored in the previous section. Indeed, for the cosmological constant or a Maxwell field, one has ${\cal T}^{t}{}_{t}={\cal T}^{r}{}_{r}$, corresponding to $\eta=\infty$, implying that the effective Hamiltonian should not depend on the field $a$ but solely on $b$. This is indeed consistent with what was derived earlier in (27) and (48). As for the massless scalar field, one has ${\cal T}^{t}{}_{t}={\cal T}^{r}{}_{r}$, indicating $\eta=1$ and thus implying that the effective Hamiltonian should be a function of c=b(1a/r)c=b(1-a/r). This is perfectly consistent with our earlier derivation of the effective Hamiltonian (34) induced by a massless scalar field, which we recall here:
@@ V.1 Polynomial ansatz in powers of mass
[…]
@@ V.2 Effective Hamiltonians for Fluid Matter
To illustrate the scope of the presented framework, let us construct the effective Hamiltonian for self-gravitating isotropic fluids, whose effective action is difficult to construct directly from a fundamental matter Lagrangian. [^5] In Ref.Brown:1992kc , an effective action of perfect fluid is proposed. Adding boundary terms properly and reducing it to spherically-symmetric static cases, the effective Hamiltonian becomes $b\rho$, although $\rho$ is not a function of $a$ and $r$ generically.
@@
Let us look for an effective Hamiltonian of type $(+1)$, thus an effective energy density $\rho_{\mathrm{eff}}[a,r]$, that fits with this evolution $(a(r),b(r))$. Let us assume, for simplicity, a quadratic truncation with
@@
$2G\rho_{\mathrm{eff}}=\frac{9}{14}+\frac{9}{98}c_{2}r^{2}-\left(\frac{1}{2r}+\frac{3}{7}c_{2}r\right)a+\frac{1}{2}c_{2}a^{2}\,$, (77)
where $c_{2}(r)$ is an arbitrary function. Let’s not forget (from (60)) that we still need to impose $\frac{\partial\rho_{\mathrm{eff}}(a,r)}{\partial a}<0$ to respect the positive energy condition. This constraint is automatically satisfied for the simplest choice $c_{2}(r)=0$, which reads: [^6] Instead of setting $c_{2}(r)=0$, noting $\eta=3/2$ for conformal fluids, we can ask the consistency between the ansatz (65) and the condition (64) up to the linear level in $a$ and $b$ and find the consistent form of $c_{2}(r)$ for a large $r$.
@@
[…]
@@
where f(r)=1a(r)/rf(r)=1-{a(r)}/{r} is the usual metric factor entering the spherically-symmetric ansatz (3). We do not have a simple argument for obtaining this surprisingly simple effective Hamiltonian. Nevertheless, we point out that $bf$ is simply $\sqrt{{-g_{tt}}/{g_{rr}}}$, that is, the speed of light for radial light-like geodesics. This might be more that a mere coincidence for the spherically-symmetric causal-limit fluid, which has no characteristic properties except for the limiting behavior $\rho_{3d}=p_{3d}$, but we postpone the investigation of a deeper origin of this effective Hamiltonian to future work. [^7] For causal-limit fluids, we have $\eta=1$ and $c=bf$ in (64). Therefore, the simple form (86) is automatically consistent with the condition (64). This could be a possible origin.